OpenStudy (anonymous):

f(x)=0.1x^4-3.2x+7 Use derivatives to find the x-coordinates of all critical points I got 0=0.4(x^3-8) I know that means 0=0.4 and 0=x^3-8. Does that mean 0=0.4 is undefined?

OpenStudy (campbell_st):

f'(x) = 0.4x^3 - 3.2 let it = 0 0.4x63 -3.2 = 0 or 0.4x^3 = 3.2 x^3 = 8 then the stationary points are at x = 2 thats the only one f"(x) = 1.2x^2 test x= 2 f"(x) >0 therefore at x = 2 there is a minimum

OpenStudy (campbell_st):

I think you'll find there is a point of inflexion at x = 0... found by solving the 2nd derivative

OpenStudy (anonymous):

Thank you!!!! :)